OpenStudy (anonymous):

If the radius of a melting snowball decreases at a rate of 1 ins/min, find the rate at which the volume is decreasing when the snowball has diameter 4 inches.

OpenStudy (ybarrap):

$$ V=(4/3 )\pi r^3\\ \cfrac{dV}{dt}=(4/3 )\pi 3r^2\times \cfrac{dr}{dt}\\ $$ $$ \cfrac{dr}{dt}=-\cfrac{1'}{min} $$ So $$ \cfrac{dV}{dt}=-(4/3 )\pi 3(4/2)^2\times 1=?\\ $$ In cubic inches per min.

OpenStudy (ybarrap):

Inches, not feet: $$ \cfrac{dr}{dt}=-\cfrac{1"}{min} $$

OpenStudy (skyz):

The Sky >.> Smiles upon you

OpenStudy (anonymous):

well thank you! @Skyz

OpenStudy (skyz):

:P I know I know nothing of the relevance to the question! But had to.